For i = the square root of -1, if 3i (2 + 5i) = x + 6i, then x = ?
A. –15
B. 5
C. 5i
D. 15i
E. 27i

Found 2 solutions by stanbon, josgarithmetic:
Answer by stanbon(75887)

For i = the square root of -1,
if 3i (2 + 5i) = x + 6i, then x = ?
-------
6i - 15 = x + 6i
----------
x = -15
------------------
Cheers,
Stan H.


A. –15
B. 5
C. 5i
D. 15i
E. 27i

Answer by josgarithmetic(39618)